# Breaking Barriers or Building Dependency? Exploring Team-LLM Collaboration in AI-infused Classroom Debate

### v1

@@ 3.3. Data Collection and Analysis
All the debate sessions were recorded, with a teaching assistant responsible for data collection and note-taking. The participants who gave consent were included in the recordings. The teaching assistant also provided operational support, such as clarifying debate rules or responding to technical issues. After debates, semi-structured interviews were conducted with all participants individually, and they were encouraged to discuss their personal experiences. The structured questions included how students interacted with the AI tool and each other (including teammates and opponents) and how they experienced these interactions. Based on their responses, more specific questions might be asked of them. The average interview duration for each participant was 30.75 minutes. The total duration of audio data collected in this study was 1022.15 minutes, encompassing recordings from all participating students. All audio data were transcribed verbatim and manually proofread, totaling 370,551 words.
First, two coders independently performed the initial coding of the data, followed by a review by a third researcher to ensure consistency and completeness. In cases of disagreement during the coding process, discussions were held to reach a consensus. For the thematic analysis (Braun and Clarke, 2012), researchers annotated the transcribed text based on conversation recordings to identify key terms, critical concepts, and collaborative behaviors. Audio recordings were instrumental in confirming certain behaviors and emotional responses, allowing for a deeper understanding of the behavioral patterns observed during the debates. Subsequently, the data were coded and clustered into hierarchical themes, through iterative analysis and discussion, we distilled 10 primary themes and 29 sub-themes that reflect participants’ shared experiences and unique perspectives. Each theme is supported by multiple representative quotes to ensure the results’ representativeness, showcasing the broader contextual significance of the sub-themes. To gain broader insights, we incorporated classroom recordings and dialogue transcripts between groups and ChatGPT as supplementary materials. These materials provided contextual examples to support the thematic analysis results, enhancing the understanding of the findings in specific contexts and validating and refining certain themes and patterns.
In three rounds of debate, participants sent 59 messages, totaling 8,181 words, and an equal number of GPT responses, totaling 37,752 words. On average, each user message contains 139 words (SD = 322), while each GPT response averages 640 words (SD = 314). Detailed statistics are shown in the Appendix B.
@@ 4.1. Ways Learners Pose Questions to AI
During the debate, when users collaborate with AI in a team setting, members ask questions differently depending on the situation, seeking more targeted answers. This study identified three distinct approaches to questioning AI.
@@ 4.1.2. Approach 2: Learners provide background information to AI to help it form appropriate responses
P(11) noted that when discussing the topic ”ornament is a crime”, he would ”first consider different scenarios, such as the periods before and after World War II, where perspectives would differ. After World War II, modernist ornament became more minimalist and utilitarian. However, if we go back to the 1920s, during the Art Deco era before the Great Depression, this viewpoint might have been seen as unconventional.” He organized the background information and input into AI. Additionally, participants provided character background information, allowing the AI to respond in a given context. P(21) explained, ”For a particular historical figure I’m interested in, I can have the AI role-play, simulating conversations with that character.” By offering detailed historical contexts, scenario settings, and character backgrounds, participants guided the AI to simulate viewpoints and attitudes from specific periods or figures more accurately, enhancing the realism and persuasiveness of the AI’s responses.
@@ 4.1.3. Approach 3: Learners provide opposing viewpoints to AI and ask for rebuttal strategies
4.2. Ways Learner Utilizes AI-Generated Content
When faced with AI-generated content, debate participants adopt various methods to filter, discuss, and refine the information to improve their own viewpoints and build consensus among group members. This section elaborates on four primary ways in which AI-generated responses are utilized. This study identified four personal filtering behaviors when analyzing participants’ interactions with AI-generated content.
@@ 4.2.4. Mode 4: Learner asks AI to explain the filtered content further
4.3. Ways Group Interacts with AI-Generated Content
@@ 4.4. Ways Learners Divide Labor in AI Usage within a Group
In our study during the debate, we observed that group members developed various roles to utilize AI effectively. Some members became AI users, who mediate between the group and the AI. This mediation involves directly reading the AI’s responses or sharing them via team communication tools such as WeChat groups. Most members tended to act as information gatherers, sharing the information they collected by posting it in group chats or verbally presenting their thoughts. These members focused on critical thinking without interacting directly with the AI. Some members opted to serve as content evaluators after receiving AI-generated responses. Their role involved filtering and organizing the responses to enhance efficiency and minimize redundancy or confusion. Some members became ad-hoc taskers, flexibly taking on specific tasks as required by the group’s needs. Group members gradually developed a collaborative mode that best suited their team dynamics. Below is a summary of these roles along with specific examples.
@@ 4.4.1. Role 1: AI user
Additionally, several participants noted that the AI user is often the focal point of the group. As P(20) remarked, ”Team members start to rely on me more,” while P(21) observed that the AI user exhibited more active participation, especially during the debate, where they frequently stood to speak. P(17) said, ”I feel like they know the most in the group because they know the AI’s responses, everyone’s arguments, and how everything is integrated; they also provide feedback to the AI.” This indicates that AI users tend to become key figures in the debate.
@@ 4.4.4. Role 4: Ad-hoc tasker
Lastly, the group’s ad-hoc tasker involves members taking on specific tasks as needed, especially in time-sensitive situations. P(6) said, ”In the early stages of research, I sometimes find an important point. If someone brings it up during the debate, I’ll respond immediately.” This flexibility allows them to respond effectively to opposing arguments, using information gathered during counter-arguments preparation. P(21) noted that during the questioning phase, when time is limited, ”one person operates the computer and talks to ChatGPT, another summarizes and extracts keywords, while the remaining three take on roles for summarizing, rebuttal, and offering supplementary explanations during the free debate phase. Each person can stand and speak based on their own views, flexibly responding to new ideas in the discussion.” This shows that every group member has a clear role and can efficiently handle information and respond to debate points.
@@ 4.5.1. Approach 1: Learners use social media to search and exchange information
First, our study found that group members frequently use social media platforms like WeChat to exchange information and platforms like Zhihu or Xiaohongshu to search for information. For instance, P(20) noted, ”I copied and pasted all AI responses into the WeChat group, while other members used different search engines, like Zhihu and Xiaohongshu, to find relevant information. We set up a WeChat group and shared all the materials we found.” P(17) added that she would ”take screenshots of collected information and then post it in the WeChat group.” Some participants also use shared mind-mapping tools for brainstorming. For example, P(12) stated, ”We used tools like mind maps in Mubu to organize arguments. I would copy AI responses and material found by other members from Zhihu or Xiaohongshu and organize them under the corresponding argument. This way, when I need to speak, I can directly extract and present the information.”
@@ 5.1. Advantage 1: Reducing Social Anxiety and Communication Barriers Among Learners
We observed that AI’s involvement significantly alleviates participants’ social anxiety in collaborative environments. In team collaboration, AI helps break discussion deadlocks and facilitates communication among group members. Specifically, AI serves as a conversational partner through a convenient, low-pressure interaction mode, reducing users’ social anxiety:
@@ 5.1.2. Reducing the Effort of Adjusting Language Based on the Person
Another way AI eases communication is by eliminating the need to tailor language based on the characteristics of different individuals. P(16) mentioned that when faced with real people, his communication is influenced by the other person’s personality, and he must adjust his language accordingly. In contrast, with AI, ”there are fewer concerns.” P(1) added, ”I feel that communicating with AI is more direct. I can ask whatever I want without considering the other person’s personality. I might adjust my questions based on their character if I were talking to a real person.” Lastly, P(20) mentioned that when AI gives an incorrect answer, he can directly ask AI to modify its response.
@@ 5.1.3. Allowing Asking Unlimited Questions Without Restrictions
Participants can ask AI questions as often as needed without feeling limited. P(2) noted, ”When communicating with people, you may have concerns, like worrying if your question seems silly. However, with AI, you can ask anything freely, from wild ideas to any other question, without worrying about such concerns.” P(16) explained that when communicating with people if your question isn’t clear the first time, they may keep asking follow-up questions until you clarify it completely. But with AI, ”you can ask questions anytime and receive immediate responses without worrying about unclear expressions.”
@@ 5.1.4. Breaking Communication Deadlocks and Facilitating Group Discussions
AI can also help group members focus on idea generation through its quick language organization. P(14) noted, ”Another advantage is that AI’s word choice is always on point. Sometimes we struggle to find the right words or use placeholders, but AI’s words are accurate and reliable.” P(6) mentioned that AI helped them rewrite a text, which ”unblocked our earlier mental blockages, allowing us to naturally continue the discussion.” These behaviors show that when interacting with AI, people can reduce their social anxiety and enjoy greater freedom of expression. Communication becomes more direct and efficient by not considering the personality, background, number of questions, or other interpersonal factors. This improves communication efficiency and allows people to express their thoughts more freely without fear of being misunderstood or offending others. In this way, AI promotes smoother information exchange.
@@ 5.2. Advantage 2: Lowering the Threshold and Providing Scaffolding for Novices
For novices, AI lowers the entry barrier to debate by quickly constructing a debate framework in a short time. It can provide a large amount of content for beginners to reference, helping them think from multiple angles about the same issue. Novices can also quickly and accurately search for and understand terminology and background information, helping them build knowledge in relevant fields. Additionally, AI can generate more professional suggestions, which are particularly valuable for less experienced users, helping them express and participate more effectively in professional discussions. AI also helps organize and clarify users’ original ideas, allowing for better structure and expression of their thoughts. Below, we explain these findings in more detail.
@@ 5.2.1. Providing Novices with Much Information Quickly
First, AI helps novices construct thought and debate frameworks by quickly providing information. For example, P(5) mentioned, ”At first, we were all quite unclear about the debate process. We only knew we had to present arguments and then engage in free debate, but we didn’t have a clear concept of how to structure an argument.” By using AI, ”the entire process is listed for you. It tells you what the opposition might say and what you need to say, forming an argument framework.” P(4) stated, ”During debate preparation, I found AI extremely helpful in organizing language quickly. It not only provided a clear framework, such as how to speak and what strategy to use, but also listed steps like introduction, explanation, and summary. The structure was very clear, and I could quickly understand and use these tips, making the preparation smoother.” P(6) added, ”Since we didn’t have much time, we couldn’t spend days thinking about arguments and crafting polished statements. AI gave us an outline, and we selected the most suitable content, greatly reducing our preparation time.”
@@ 5.2.2. Helping Novices Better Organize and Express Their Ideas
In debates, novice participants mentioned that AI helped them organize and structure their thoughts. For instance, P(3) said, ”I gave AI some of my ideas or a long description without needing to summarize it myself. Sometimes it was just a simple description of an objective fact or a collection of random thoughts, and AI extracted the key points and organized them.” P(7) added, ”We encountered problems that we couldn’t structure with our current knowledge. In this case, we used AI. AI gave us a logical framework, which became the foundation for clearer discussions.” Participants also noted that AI helps them quickly generate language that conforms to debate norms. P(2) said, ”I’m not very good at speaking in public, but with AI’s help, I now know how to start, like ’Dear judges, esteemed opponents… First, our side believes…’ I didn’t know these before, but AI quickly helped me structure my points in a specific format. ”
@@
During debates, AI provides a large amount of information that requires participants to spend time processing, leading to information overload. Due to the high intensity of debates, participants’ cognitive resources become particularly valuable, making them more sensitive to information overload. In such high-pressure environments, excessive or irrelevant information can disrupt participants’ train of thought.
@@ 5.3.1. Providing Too Many Ideas
First, participants mentioned that AI provided too many ideas, which increased the time needed for understanding. For example, P(8) stated, ”I feel like AI can generate a lot of ideas in a short period, but first, we need time to process and understand, which leaves me with no time for personal thinking because AI is too efficient.” Participants also noted that during short scenarios, such as cross-examinations, ”I first had to understand the content provided by AI, then supplement what was missing, which took quite a bit of time.” This demonstrates how an overabundance of AI-generated answers can lead to cognitive load (Sweller, 2011). Users may struggle to fully develop and explore their own ideas when they are busy understanding and improving AI-generated thoughts.
@@ 5.4.2. Reducing Individual Autonomy
Some participants felt that the AI-determined arguments in the debate process lowered their personal involvement. P(1) stated, ”The arguments didn’t come from me.” Participants may not feel strong ownership of these AI-generated arguments. P(8) mentioned that AI reduced the ”sense of achievement that comes from going from zero to one.” P(6) echoed, ”The concluding and opening speeches were almost entirely based on AI’s content. While convenient and quick, it lacked personal input, and felt a bit disappointed,”
Additionally, participants felt that interacting with AI lacked the stimulating and engaging experience of real human interaction. P(5) emphasized that AI lacks the interactive dialogue experience and emotional connection that make conversations enjoyable and meaningful. He felt that AI could not be considered a true debate participant but an auxiliary tool. He said that using AI ”doesn’t reflect the true meaning of debate” since it should represent personal thoughts and opinions, which made him feel guilty.
@@ 5.5.1. Incomplete and Inconsistent in AI-Generated Information
Participants noted that AI’s short memory for text caused conversation content to be incomplete and inconsistent. P(8) mentioned, ”AI doesn’t have a long-term memory. For example, if I ask it to respond in Chinese, after a few minutes, it switches back to English.” P(18) similarly pointed out, ”AI doesn’t absorb everything mentioned in a conversation. The content we discussed earlier often gets forgotten after a few rounds.” He also pointed out, ”Sometimes you ask AI for something very specific, like ’Who said this quote?’ or ’Which era is this from?’ AI sometimes gives wrong answers. So during debates, we didn’t ask it much because the information it gave wasn’t useful,” P(11) explained. When AI made such mistakes, participants expressed a decline in trust. P(21) said, ”Later on, we didn’t trust AI as much. We relied more on our own thinking and stopped using AI as much.”

### v2

@@ 3.3. Data Collection and Analysis
All the debate sessions were recorded, with a teaching assistant responsible for data collection and note-taking. The participants who gave consent were included in the recordings. The teaching assistant also provided operational support, such as clarifying debate rules or responding to technical issues. After debates, semi-structured interviews were conducted with all participants individually, and they were encouraged to discuss their personal experiences. The structured questions included how students interacted with the AI tool and each other (including teammates and opponents) and how they experienced these interactions. Based on their responses, more specific and detailed questions might be asked. The average interview duration for each participant was 30.75 minutes. The total duration of audio data collected in this study was 1022.15 minutes, including recordings from all participating students. All audio data were transcribed and manually proofread, totaling 370,551 words.
First, two coders independently coded the data’s initial coding, followed by a review by a third researcher to ensure consistency and completeness. In cases of disagreement during the coding process, discussions were held to reach a consensus. For the thematic analysis (Braun and Clarke, 2012), researchers annotated the transcribed text based on audio recordings to identify key terms, critical concepts, and collaborative behaviors. Audio recordings were instrumental in confirming certain behaviors and emotional responses, allowing for a deeper understanding of the behavioral patterns observed during the debates. Subsequently, the data were coded and clustered into hierarchical themes. Through iterative analysis and discussion, we distilled 10 primary themes and 29 sub-themes that reflect participants’ shared experiences and unique perspectives. Each theme is supported by multiple representative quotes to ensure the results’ representativeness, showcasing the broader contextual significance of the sub-themes. To gain broader insights, we incorporated classroom recordings and dialogue transcripts between groups and ChatGPT as supplementary materials. These materials provided contextual examples to support the thematic analysis results, enhancing the understanding of the findings in specific contexts and validating and refining certain themes and patterns.
In three debate rounds, participants sent 59 messages, totaling 8,181 words, and an equal number of GPT responses, totaling 37,752 words. On average, each user message contains 139 words (SD = 322), while each GPT response averages 640 words (SD = 314). Detailed statistics are shown in the Appendix B.
@@ 4.1. Ways Learners Pose Questions to AI
During the debate, when learners collaborate with AI in a team setting, they ask questions differently depending on the situation, seeking more targeted answers. This study identified three distinct approaches to questioning AI.
@@ 4.1.2. Approach 2: Learners provide background information to AI to help it form appropriate responses
P(11) noted that when discussing the topic ”ornament is a crime”, he would ”first consider different scenarios, such as the periods before and after World War II, where perspectives would differ. After World War II, modernist ornament became more minimalist and utilitarian. However, if we go back to the 1920s, during the Art Deco era before the Great Depression, this viewpoint might have been seen as unconventional.” He organized the background information and input into AI. Additionally, participants provided character background information, allowing the AI to respond in a given context.
P(21) explained, ”For a particular historical figure I’m interested in, I can have the AI role-play, simulating conversations with that character.” By offering detailed historical contexts, scenario settings, and character backgrounds, participants guided the AI to simulate viewpoints and attitudes from specific periods or figures more accurately, enhancing the authenticity and persuasiveness of the AI’s responses.
@@ 4.1.3. Approach 3: Learners provide opposing viewpoints to AI and ask for rebuttal strategies
4.2. Ways Individual Learner Utilizes AI-Generated Content
When faced with AI-generated content, debate participants adopt various methods to filter, discuss, and refine the information to improve their own viewpoints and build consensus among group members. This section elaborates on four primary ways in which AI-generated responses are utilized. This study identified four individual filtering behaviors when analyzing participants’ interactions with AI-generated content.
@@ 4.2.4. Mode 4: Learner asks AI to explain the filtered content further
4.3. Ways Group Members Interact with AI-Generated Content
@@ 4.4. Ways Learners Divide Labor in AI Usage within a Group
In our study during the debate, we observed that group members developed various roles to utilize AI effectively. Some members became AI users, who mediate between the group and the AI. This mediation involves directly reading the AI’s responses or sharing them via communication apps like WeChat. Most members tended to act as information gatherers, sharing the information they collected by posting it in group chats or verbally presenting their thoughts. These members focused on critical thinking without interacting directly with the AI. Some members opted to serve as content evaluators after receiving AI-generated responses. Their role involved filtering and organizing the responses to enhance efficiency and minimize redundancy or confusion. Some members became ad-hoc taskers, flexibly taking on tasks as required by the group’s needs. Group members gradually developed a collaborative mode that best suited their team dynamics. Below is a summary of these roles, along with examples.
@@ 4.4.1. Role 1: AI user
Additionally, several participants noted that the AI user is often the group’s focal point. As P(20) remarked, ”Team members start to rely on me more,” while P(21) observed that the AI user exhibited more active participation, especially during the debate, where they frequently stood to speak. P(17) said, ”I feel like they know the most in the group because they know the AI’s responses, everyone’s arguments, and how everything is integrated; they also provide feedback to the AI.” This indicates that AI users tend to become key figures in the debate.
@@ 4.4.4. Role 4: Ad-hoc tasker
Lastly, the group’s ad-hoc tasker involves members taking on specific tasks as needed, especially in time-sensitive situations. P(6) said, ”In the early stages of research, I sometimes find an important point. I’ll respond immediately if someone brings it up during the debate.” This flexibility allows them to respond effectively to opposing arguments, using information gathered during counter-arguments preparation. P(21) noted that during the questioning phase, when time is limited, ”one person operates the computer and talks to ChatGPT, another summarizes and extracts keywords, while the remaining three take on roles for summarizing, rebuttal, and offering supplementary explanations during the free debate phase. Each person can stand and speak based on their own views, flexibly responding to new ideas in the discussion.” This shows that every group member has a clear role and can efficiently handle information and respond to debate points.
@@ 4.5.1. Approach 1: Learners use social media to search and exchange information
First, our study found that group members frequently use communication apps like WeChat to exchange information and social media like Zhihu or RedNote to search for information. For instance, P(20) noted, ”I copied and pasted all AI responses into the WeChat group, while other members used different search engines, like Zhihu and RedNote, to find relevant information. We set up a WeChat group and shared all the materials we found.” P(17) added that she would ”take screenshots of collected information and then post it in the WeChat group.” Some participants also use shared mind-mapping apps for brainstorming. For example, P(12) stated, ”We used tools like mind maps in Mubu to organize arguments. I would copy AI responses and material found by other members from Zhihu or RedNote and organize them under the corresponding argument. This way, when I need to speak, I can directly extract and present the information.”
@@ 5.1. Advantage 1: Reducing Social Anxiety and Communication Barriers Among Learners
We observed that AI’s involvement significantly alleviates participants’ social anxiety in collaborative environments. In team collaboration, AI helps break discussion deadlocks and facilitates communication among group members. Specifically, AI serves as a conversational partner through a convenient, low-pressure interaction mode, reducing users’ social anxiety.
@@ 5.1.2. Reducing the Effort of Adjusting Language Based on the Person
Another way AI eases communication is by eliminating the need to tailor language based on the characteristics of different individuals. P(16) mentioned that when faced with real people, his communication is influenced by the other person’s personality, and he must adjust his language accordingly. In contrast, with AI, ”there are fewer concerns.” P(1) added, ”I feel that communicating with AI is more direct. I can ask whatever I want without considering the other person’s personality. If I were talking to a real person, I might adjust my questions based on their character.” Lastly, P(20) mentioned that when AI gives an incorrect answer, he can directly ask AI to modify its response.
@@ 5.1.3. Allowing Asking Unlimited Questions Without Restrictions
Participants can ask AI questions as often as needed without feeling limited. P(2) noted, ”When communicating with people, you may have concerns, like worrying if your question seems silly. However, with AI, you can ask anything freely, from wild ideas to any other question, without worrying about such concerns.” P(16) explained that when communicating with people, if your question isn’t clear the first time, they may keep asking follow-up questions until you clarify it completely. But with AI, ”you can ask questions anytime and receive immediate responses without worrying about unclear expressions.”
@@ 5.1.4. Breaking Communication Deadlocks and Facilitating Group Discussions
AI can also help group members focus on idea generation through its quick language organization. P(14) noted, ”Another advantage is that AI’s word choice is always on point. Sometimes we struggle to find the right words or use placeholders, but AI’s words are accurate and reliable.” P(6) mentioned that AI helped them rewrite the text, which ”unblocked our earlier mental blockages, allowing us to naturally continue the discussion.”
These behaviors show that when interacting with AI, learners can reduce their social anxiety and enjoy greater freedom of expression. Communication becomes more direct and efficient by not considering the personality, background, number of questions, or other interpersonal factors. This improves communication efficiency and allows people to express their thoughts more freely without fear of being misunderstood or offending others. In this way, AI promotes smoother information exchange.
@@ 5.2. Advantage 2: Lowering the Threshold and Providing Scaffolding for Novices
For novices, AI lowers the entry barrier to debate by quickly constructing a debate framework in a short time. It can provide a large amount of content for beginners to reference, helping them think from multiple angles about the same issue. Novices can also quickly and accurately search for and understand terminology and background information, helping them build knowledge in relevant fields. Additionally, AI can generate professional suggestions, particularly valuable for less experienced users, helping them express and participate more effectively in professional discussions. AI also helps organize and clarify learners’ original ideas, allowing for better structure and expression of their thoughts. Below, we explain these findings in more detail.
@@ 5.2.1. Providing Novices with Much Information Quickly
First, AI helps novices construct thought and debate frameworks by quickly providing information. For example, P(5) mentioned, ”At first, we were all quite unclear about the debate process. We only knew we had to present arguments and then engage in free debate, but we didn’t have a clear concept of how to structure an argument.” By using AI, ”the entire process is listed for you. It tells you what the opposition might say and what you need to say, forming an argument framework.” P(4) stated, ”During debate preparation, I found AI extremely helpful in organizing language quickly. It provided a clear framework, such as how to speak and what strategy to use, and also listed steps like introduction, explanation, and summary. The structure was very clear, and I could quickly understand and use these tips, making the preparation smoother.” P(6) added, ”Since we didn’t have much time, we couldn’t spend days thinking about arguments and crafting polished statements. AI gave us an outline, and we selected the most suitable content, greatly reducing our preparation time.”
@@ 5.2.2. Helping Novices Better Organize and Express Their Ideas
In debates, novice participants mentioned that AI helped them organize and structure their thoughts. For instance, P(3) said, ”I gave AI some of my ideas or a long description without needing to summarize it myself. Sometimes, it was just a simple description of an objective fact or a collection of random thoughts, and AI extracted the key points and organized them.” P(7) added, ”We encountered problems that we couldn’t structure with our current knowledge. In this case, we used AI. AI gave us a logical framework, which became the foundation for clearer discussions.” Participants also noted that AI helps them quickly generate language that conforms to debate norms. P(2) said, ”I’m not very good at speaking in public, but with AI’s help, I now know how to start, like ’Dear judges, esteemed opponents… First, our side believes…’ I didn’t know these before, but AI quickly helped me structure my points in a specific format.”
@@
During debates, AI provides a large amount of information that requires participants to spend time processing, leading to information overload. Due to the high intensity of debates, participants’ cognitive resources become particularly valuable, making them more sensitive to information overload. Excessive or irrelevant information can disrupt participants’ train of thought in such high-pressure environments.
@@ 5.3.1. Providing Too Many Ideas
First, participants mentioned that AI provided too many ideas, which increased the time needed for understanding. For example, P(8) stated, ”I feel like AI can generate a lot of ideas in a short period, but first, we need time to process and understand, which leaves me with no time for personal thinking because AI is too efficient.” Participants also noted that during short scenarios, such as cross-examinations, ”I first had to understand the content provided by AI, then supplement what was missing, which took quite a bit of time.” This demonstrates how an overabundance of AI-generated answers can lead to cognitive load (Sweller, 2011). Users may struggle to fully develop and explore their own ideas when they are busy understanding and improving AI-generated content.
@@ 5.4.2. Reducing Individual Autonomy
Some participants felt that the AI-determined arguments in the debate process lowered their feelings of involvement. P(1) stated, ”The arguments didn’t come from me.” Participants may not feel strong ownership of these AI-generated arguments. P(8) mentioned that AI reduced the ”sense of achievement that comes from going from zero to one.” P(6) echoed, ”The concluding and opening speeches were almost entirely based on AI’s content. While convenient and quick, it lacked personal input, and felt a bit disappointed,”
Additionally, participants felt that interacting with AI lacked real human interaction’s stimulating and engaging experience. P(5) emphasized that AI lacks the interactive dialogue experience and emotional connection that make conversations enjoyable and meaningful. He felt that AI could not be considered a true debate participant but an auxiliary tool. He said that using AI ”doesn’t reflect the true meaning of debate” since it should represent personal thoughts and opinions, which made him feel guilty.
@@ 5.5.1. Incomplete and Inconsistent in AI-Generated Information
Participants noted that AI’s limited text remembering and content alignment capability caused conversation content to be incomplete and inconsistent. P(8) mentioned, ”AI doesn’t have a long-term memory. For example, if I ask it to respond in Chinese, after a few minutes, it switches back to English.” P(18) similarly pointed out, ”AI doesn’t absorb everything mentioned in a conversation. The content we discussed earlier often gets forgotten after a few rounds.” He also pointed out, ”Sometimes you ask AI for something very specific, like ’Who said this quote?’ or ’Which era is this from?’ AI sometimes gives wrong answers. So during debates, we didn’t ask it much because the information it gave wasn’t useful,” P(11) explained. When AI made such mistakes, participants expressed a decline in trust. P(21) said, ”Later on, we didn’t trust AI as much. We relied more on our own thinking and stopped using AI as much.”